Bella18:

Ok so I'm on this math problem and I need help its 2/16× × =18/32

Tbone:

So you're going to want to do 2 over 16(18 over 32 heres how I did it. But I think you have to simplify

Bella18:

Ok ty that helped

idfkaa:

@tbone wrote:
So you're going to want to do 2 over 16(18 over 32 heres how I did it. But I think you have to simplify
damn tbone smart

idfkaa:

@bella18 wrote:
Ok ty that helped
is this algebra 2

Bella18:

Yes

umm:

I don't feel like Tbone's response is what you were looking for... So I'll try my best to answer this.

umm:

Given: \[\frac{ 2 }{ 16 } \times x\] \[x = \frac{ 18 }{ 32 }\] It would look like this: \[\frac{ 2 }{ 16 }\left(\begin{matrix}18 \\ 32\end{matrix}\right)\] Now, if you apply the rule of \[(a)=a\] It would look like this: \[\left(\begin{matrix}18 \\ 32\end{matrix}\right) = \frac{ 18 }{ 32 }\] So therefore for part 1, it'd completely look like this:\[= \frac{ 2 }{ 16 } \times \frac{ 18 }{ 32 }\] Now, you'd just need to cancel the numbers: \[\frac{ 2 }{ 16 } = \frac{ 1 }{ 8 }\] \[\frac{ 1 }{ 8 } \times \frac{ 18 }{ 32 }\] In order to do that, we use factoring, which will look like this on the second fraction: \[\frac{ 18 }{ 32 } = \frac{ 9 }{ 16 }\] Factoring: \[18=2\times9\] \[=\frac{ 2\times9 }{ 32 }\] Factoring continued:\[32=2\times 16\] \[=\frac{ 2\times9 }{ 2\times16 }\] *Canceling the common factor: 2 \[=\frac{ 9 }{ 16 }\]\[=\frac{ 1 }{ 8 }\times \frac{ 9 }{ 16 }\] Applying the fraction rule; example: \[\frac{ a }{ b} \times \frac{ c }{ d }=\frac{ a \times c }{ b \times d }\] With our problem: \[=\frac{ 1 \times 9 }{ 8 \times 16}\] Multiplication: \[1 \times 9 = 9\] \[8 \times 16 = 128\] Which leaves you with your final answer: \[=\frac{ 9 }{ 128 }\] *For such terms as canceling for fractions, it's basically dividing a smaller number for both the top number and bottom number which is our common factor. Here's an example \[\frac{ 3 }{ 9 } = \frac{ 1 }{ 3 }\] Which leads 3 to be your common factor. Hope this has helped, and good luck! (: